### v1

# Exploring Concept Depth: How Large Language Models Acquire Knowledge at Different Layers?

Abstract
This paper studies the phenomenon that different concepts are learned in different layers of large language models, i.e. more difficult concepts are fully acquired with deeper layers. We define the difficulty of concepts by the level of abstraction, and here it is crudely categorized by factual, emotional, and inferential. Each category contains a spectrum of tasks, arranged from simple to complex. For example, within the factual dimension, tasks range from lie detection to categorizing mathematical problems. We employ a probing technique to extract representations from different layers of the model and apply these to classification tasks. Our findings reveal that models tend to efficiently classify simpler tasks, indicating that these concepts are learned in shallower layers. Conversely, more complex tasks may only be discernible at deeper layers, if at all. This paper explores the implications of these findings for our understanding of model learning processes and internal representations. Our implementation is available at https://github.com/Luckfort/CD.
[…]
123
1 Introduction
Modern large language models (LLMs) such as GPT-4 [1], Gemma [29], Llama [32] have impressive generation and inference capabilities. It is a common insight that large models encode a large amount of knowledge within their parameters and that a larger number of parameters leads to better performance. Consequently, popular large models are usually very deep networks containing billions or even trillions of parameters [16, 13]. It is observed that as the size of the model parameters increases, the model exhibits dramatic and unpredictable performance improvements on specific tasks, which are so-called emergent abilities [33]. For instance, GPT-3 [6] shows a large increase in performance after exceeding 13B parameter in size, and a similar phenomenon was observed for LaMDA [31] after exceeding 68B parameter [34]. However, it remains obscure how and where LLMs actually understand the concept of knowledge correctly, and there is currently a lack of empirically summarized research on the relationship between LLMs’ network depth and conceptual understanding.
Recent studies usually approach concerns above from the following two perspectives, 1) analyzing model weights and architectures [18], e.g., the pruning method removes weights based on their importance and reverse inference of where the model learns the concept of knowledge by determining whether it will affect the model performance. No matter it is single pruning [18] or iterative pruning [15], existing pruning methods applied to LLMs are usually designed to be relatively complex and not generalizable, requiring adjustments for different datasets [23]. 2) probing representations [2], which typically utilize linear classifier probes to test the performance of the hidden layer representations. For example, the most recent probing-based work [16] explores each layer’s performance of the Llama series on the counterfactual dataset [24] with representations from different layers.
In this paper, we propose the notion of "Concept Depth" to measure where different models, on different types of datasets, understand knowledge. To our knowledge, this is the first time this notion has been introduced in related work. We draw empirical summaries from extensive experiments on the "Concept Depth" of existing mainstream large language models on different types of datasets through probe techniques. As shown in Figure 1, we capture the feature responses of different layers of the LLMs for different datasets and use independent linear probes to indicate the best performance that the current layer can achieve. The core contribution of our work can be summarized as follows:
Figure 1: a) The overview of process of our analysis. LLMs respond to text prompts. The probing process evaluates the optimal performance that the current layer of LLMs can achieve. b) A conceptual example of layer-wise representations for probe detection. Orange points and Blue points represent fake samples and real samples, respectively. The probing task here is the binary classification on different types of samples.
Concept Depth. Introducing the concept of "concept depth" to measure where a concept is learned by different size LLMs. Our experiments show that basic concepts are often learned at a low level of "concept depth", while more complex concepts require more depth. This is a consistent phenomenon across LLMs of different model families and different sizes.
Deconstructing LLMs Capabilities. Deconstructing LLM capabilities by analogy with the composition of complex human capabilities. We conduct experiments on a large number of relevant datasets, synthesize comprehensive performance summaries of different LLMs, and obtain specific concept depths of optimal performance for different capabilities.
Robustness from Concept Depth perspective. Providing a new perspective on LLMs robustness. We conduct ablation experiments on other factors (e.g., quantization and random noise) that may affect the accuracy of LLMs inference, which showed that these interfering factors affect the concept depth of the model so that the LLMs appear to end up learning concepts at a deeper location than before. It can be assumed that deeper models have more concept depth to counteract interference.
2 Related Work
2.1 Concepts representation in DNNs
Identifying similarities across various examples to form concepts, plays a crucial role in human decision-making [3]. Many studies have explained DNNs’ (Deep Neural Networks) decision-making based on a conceptual perspective, describing the global behavior of DNNs in a human-understandable way [35, 27, 28, 8]. For example, Yeh et al. [35] demonstrated that DNNs exhibit conceptual representations through the activation patterns observed in their hidden or output layers. Further, Räz [27] indicates that DNNs learned not only conceptual representations of predicted categories but also indirect concepts that contribute to the prediction. A notable study reveals the existence of a ‘representation bottleneck,’ highlighting a cognitive disparity between DNNs and humans. This phenomenon is characterized by DNNs’ tendency to grasp representations of concepts that are either too simple or overly complex, while they often struggle with acquiring representations of concepts of moderate complexity [8]. Motivated by previous work, our paper aims to cover concepts of different complexities to further understand concepts within LLMs.
2.2 Knowledge and concepts in LLMs
The great performance of the LLMs in various downstream tasks (e.g. LLMs can predict factual statements about the world based on prompts [24].) has led to a great discussion about whether these capabilities are ‘stochastic parrots’ or LLMs really understands these concepts. Pioneeringly, Gurnee and Tegmark [14] showed that LLMs internally store concepts like latitude, longitude, and time. Similarly, another work showed that the truth of a statement can be detected from the internal state of LLMs [4]. Geva et al. [12] also came to similar conclusions by artificially blocking or “knocking out" specific parts of the LLMs to observe their effects on the inference process. These related studies show the existence of structures for understanding concepts within LLMs, which motivates us to explore how concepts at various complexities are encoded within various depths of LLMs.
2.3 Explorations of Interpretability in LLMs
A large number of related studies have deconstructed the inner layers of LLMs from various perspectives to understand the mechanisms inside such models [36]. Fan et al. [10] computes stopping signals by evaluating key features to early stop the LLM inference and get the internal performance of the LLMs, concluding that not all layers are necessary. Through pruning the LLMs, Gromov et al. [13] found that the parameters of some layers were not utilized correctly. Men et al. [23] also shows a high level of redundancy in the LLM architecture. Probes trained with logistic regression are a well-established method [2] that has been applied in classifying the truthfulness of LLMs and has been validated in many studies [22, 4, 19]. The latest work related to us detects different layers in Llama series responding to facts or counterfactuals by probing techniques [16]. Inspired by these work, we propose Concept Depth to summarize these phenomena. Our work is different from previous work. Our work is relatively comprehensive, and we focus on the Concept Depth of where it appears in the LLMs, analyzing it experimentally by training linear classifier probes.
Figure 2: The LLMs are trying to understand easy tasks and complex tasks. The more complex the task is, the deeper layers an LLM needs to understand it. The stronger the LLM is, the more complex the task level it can learn.
3 Analyzing Method
In this paper, we designed a complete probing framework to under-stand how concepts at various levels are encoded within LLMs andinvestigate whether LLMs’ internal representations are robust withrespect to concept.
3.1 Visualizing Layer-wise Representations
Contemporary LLMs typically consist of numerous layers of deep Transformer decoders (see Table 1), yet the conceptual significance of the encoded information within each layer remains largely unexplored. To illustrate this point, we analyzed the representations from every layer across an entire LLM, employing principal component analysis (PCA) to facilitate visualization. Specifically, for each layer’s representation, we will downscale the representation to two dimensions and visualize it following the gold label (e.g., math problem or not). The comparison across layers at various depths will indicate LLMs’ understanding of problems in need of different concepts. For instance, as shown in Figure 1, we project the representation of Counterfact [24] in the 80% depth of Gemma-7b, and the label is the real answer of the data set. It can be observed that gemma-7b has significantly different responses for inputs with different type answers at this depth.
3.2 Linear Classifier Probing
To better understand how representations from different layers could contribute to the final prediction, we utilized a linear classifier to further quantify the probing differences. For one specific task $w$, the hidden feature set in LLMs is $x\in\mathbb{R}^{n\times d_{model}}$, where $n$ denotes number of samples, and $x^{(i)}\in\mathbb{R}^{1\times d_{model}}$ represent the representation at a certain layer, where $d_{model}$ donate the dimension for the hidden layer representation and the details of $d_{model}$ for each LLM can be found in the appendix. Binary label $y^{(i)}$ is set as 0 or 1. The objective function of such a binary Logistic regression classifier probes with L2 regularization can be written as:
$\scriptsize J(\theta)=-\frac{1}{n}\sum_{i=1}^{n}Cost(\sigma(x^{(i)}),y^{(i)})+\frac{\lambda}{2n}\sum_{j=1}^{m}\theta_{j}^{2}$ (1)
and
$\scriptsize Cost(\sigma(x^{(i)}),y^{(i)})=y^{(i)}\log\left(\sigma(\theta^{T}x^{(i)})\right)+(1-y^{(i)})\log\left(1-\sigma(\theta^{T}x^{(i)})\right)$ (2)
where $\theta$ is the parameter in this Logistic regression model, $\lambda$ is the regularization parameter. The linear model predicts LLM’s response to the test set, compared with the true label. This yields a quantification of how well LLMs understand the current depth.
Refer to caption (a) Refer to caption (b)
Refer to caption (c) Refer to caption (d)
Figure 3: Analysis diagrams of Section 5.2. The converge proportion and peak accuracy of each model over the nine datasets. (a) and (c) shows the converge proportion over the datasets. (b) and (d) shows the peak accuracy over the datasets.
4 Experimental Setting
Our experiments used 9 datasets containing three aspects (human emotions, logical reasoning, and factual analysis) and categorized these 9 datasets into easy and complex datasets according to the performance of the LLMs on each dataset. Specifically, datasets in which the linear probes can obtain high classification accuracy at the initial or middle depth of the LLMs are categorized as easy datasets while other datasets with large fluctuations and stable classification accuracy occurring at the deep depth of the LLMs are categorized as complex datasets. In Section 4.1, we introduce the LLMs used for experiments. And give a description of these nine datasets in Section 4.2.
4.1 Models
In this paper, we employ three families of open-source models: Gemma (2B, 7B) [30], LLaMA (7B, 13B) [32], and QWen (0.5B, 1.8B, 4B, 7B, and 14B) [5] to support our analysis. The number of layers in each model is shown in Table 1. During probing analysis, we choose a linear classifier probe in the experiments. The ratio of training set to test set is 8:2. We follow the usual approach of LLMs probe classifier [9, 26], where the feature representation of the final transformer layer at each layer of LLMs(e.g. 14-th ’post_attention_layernorm’ in Llama2-7b (32 Layers in total) is as input to the probe. We provide the detailed structure of Llama2-7b model in appendix. The other series of models follow a similar structure.
Table 1: Number of layers in each LLM.
Model Layer Model Layer Model Layer
Gemma-2B 18 Gemma-7B 28 LLaMA-7B 32
LLaMA-13B 40 QWen-0.5B 24 QWen-1.8B 24
QWen-4B 40 QWen-7B 32 QWen-14B 40
4.2 Datasets
We used nine datasets on fact (Cities[22], CommonClaim[7], Counterfact[24]), emotion (STSA[17], IMDb[20], Sarcasm[25], HateEval [21]), and inference (StrategyQA[11], Coinflip[34]) for our experiments, and sample examples of each are given in the Appendix, Section A.2, following is a brief description of these datasets. The number of samples we use for each dataset is shown at Table 2.
Table 2: Number of samples we use in each dataset.
Dataset Size Dataset Size Dataset Size
Cities 1496 CommonClaim 6000 Counterfact 4000
HateEval 6000 STSA 6920 IMDb 2000
Sarcasm 6000 StrategyQA 2290 Coinflip 500
Cities [22]: consists of statements about the location of cities and their veracity labels (e.g., The city of Zagreb is in Japan, which is wrong).
CommonClaim [7]: A dataset of boolean statements, each labeled by two humans as common-knowledge-true, common-knowledge-false, or neither.
Counterfact [24]: Counterfact includes thousands of counterfactuals along with text that allows quantitative testing of specificity and generalization when learning a counterfactual.
Figure 4: Analysis diagrams of Section 5.1. Linear probing accuracy of three LLM families (Gemma, LLaMA, QWen) on nine datasets.
Figure 5: Analysis diagrams of Section 5.3. Linear probing accuracy of Gemma-7B, LLaMA-7B, QWen-7B on nine datasets.
Figure 6: The upper radar image is the converging point of each dataset on Gemma-7B, LLaMA-7B, and QWen-7B, represented by the percent depth proportion. The bottom radar image is the maximum accuracy of each dataset on Gemma-7B, LLaMA-7B, and QWen-7B, represented by the percentage depth proportion.
HateEval [21]: HateEval has English tweets which were annotated hierarchically.
STSA [17]: STSA includes movie reviews, half of which were considered positive and the other half negative. Each label is extracted from a longer movie review and reflects the writer’s overall intention for this review.
IMDb [20]: IMDb is a benchmark dataset for binary sentiment classification.
Sarcasm [25]: Sarcasm is a high-quality news headlines dataset that is annotated as sarcastic or not sarcastic.
StrategyQA [11]: StrategyQA contains questions across all knowledge domains to elicit creative and diverse yes/no questions that require implicit reasoning steps.
Coinflip [34]: Coinflip includes questions about coin flipping. This task requires the model to determine if a coin remains heads up after it is either flipped or left unflipped by individuals.
4.3 The Robustness of Internal Representations
4.3.1 Adding Noise
To quantify the robustness of the LLM with respect to its internal representation, when we input these questions into the LLM, we add a random string of noise in front of the input question to interfere the inference prosecc. For instance, we transform a given input question $q=\{q_{1},...,q_{i}\}$ into $q^{\prime}=\{N,q_{1},...,q_{i}\}$. Here $q_{i}$ represents the single sentence in question $q$ and $N$ is the noise string. The noise, which disrupts the classification tasks, originates from two separate random strings. These strings are inserted at the beginning of the sentence with an equal probability of 50%.
4.3.2 Quantization Settings
Quantization has been a rule of thumb for faster LLM inference. The general fact is that by using lower precision level allows LLMs to be run on less capable hardware with an acceptable reduction in their ability and accuracy. Our approach to applying quantization is to explore its effect on Concept Depth, whether it still maintains the previous Concept Depth or slows down the process of understanding knowledge. Quantization is set as 8-bit, 16-bit, and full 32-bit.
4.4 Metrics for Accuracy Variation
Definition 1 (Variation Rate) Given an LLM probing classifier $M=\{q,y,z,d\}$ ($q$, $y$, $z$, and $d$ are the input question, ground truth binary label, predicted label and total amount of layers, respectively), it has the accuracy at i-th layer $\alpha_{i}=\frac{1}{|z|}*\sum_{k=1}^{|z|}{[y_{k}=z_{k}]},i\in\{0,1,2,...,d-1\}$. We denote the variation rate $\beta_{i}=\alpha_{i}/\alpha_{i-1},i\in\{1,2,...,d-1\}$.
We introduce two metrics to capture variations in accuracy: the jump point and the converging point and define them by the given the definition of variation rate.
Definition 2 (Jump Point) We denote $J(M,D)=\min\{\frac{i}{d}\}$ s.t. $\beta_{i}>=1.1$, $i\in\{1,2,...,d-1\}$, as the jump point, where M and $D=(q,y)$ represents the LLM classifier and the dataset.
When a noticeable boost in accuracy is observed, the jump point signals the model’s recognition of critical features or patterns within the dataset.
Definition 3 (Converging Point) We denote $C(M,D)=\max\{\frac{i}{d}\}$ s.t. $|\beta_{i}-1|<0.03$, $i\in\{1,2,...,d-1\}$, as the converging point, where $M$ and $D=(q,y)$ represents the LLM classifier and the dataset.
As the accuracy plateaus or begins to decline, the convergence point indicates the model’s learning saturation or its peak learning capacity from the dataset.
Analyzing these metrics provides deeper insight into the model’s learning dynamics and its adaptability to various data types.
5 Experimental Analysis
We conduct experiments to answer the following research questions about the concept depth:
RQ1: Do different LLMs’ concept depths behave consistently in the same dataset? (Section 5.1)
RQ2: Do different size LLMs but the same series (e.g., Gemma series) have consistent Concept Depth? (Section 5.2)
RQ3: Do LLMs’ Concept Depth of the same size behave consistently? (Section 5.3)
5.1 Comparison Among the Datasets
In this section, we delve into an evaluative comparison of performance across a range of datasets, utilizing Figure 4 to detail the layer-wise accuracy of all nine LLMs over nine distinct datasets. Table 3 provides the accuracy, F1-score, and AUC metrics for parts of the layers (first, 25% depth, 50% depth, 67% depth, 83% depth, last) of each model over the nine datasets from the quantitative side. A performance threshold of 0.7 accuracy is applied to assess the models’ effective comprehension of concepts. This examination leads to two general observations. Firstly, regarding different concepts, LLMs exhibit varying accuracy trends across their layers; within individual concepts, however, different LLMs tend to display consistent accuracy patterns across these layers. Secondly, in tasks that require varying levels of conceptual understanding, the LLMs demonstrate their understanding across different neural network layers, indicating a layered approach to processing complex concepts.
Among the three factual concept datasets, significant variations in trends are observed across the models. The Cities dataset exhibits a sharp increase in comprehension at the lower layers, stabilizing in the higher layers, indicating a strong grasp of the concept. The CommonClaim dataset has become stable in early layers. On the other hand, the accuracy improvement of the nine LLMs trained on Counterfact was relatively difficult to achieve, utilizing deeper layers, and the accuracy was lower than that of many other datasets. Therefore, we can conclude that Counterfact is a complex task compared to the other datasets.
In datasets centered on emotional concept comprehension (STSA, IMDb, Sarcasm, and HateEval), despite varying levels of understanding, all models demonstrate a rise in accuracy at the initial layers, with convergence occurring in the intermediate layers. Although HateEval essentially reaches stability in the second or third layer, its accuracy reaches up to 0.8. This pattern suggests that LLMs primarily aggregate representations from lower layers to grasp emotional concepts.
Meanwhile, the datasets StrategyQA and Coinflip, which demand specific reasoning skills, tend to display a bell-shaped accuracy trajectory in all models, with peak accuracy observed in the middle layers. Such patterns underscore the intricate complexity associated with reasoning tasks.
The performance of classification tasks can be categorized into three types. 1) For Cities, STSA, IMDb, and Sarcasm, the LLMs suddenly understand the tasks at intermediate layers. 2) For CommonClaim and HateEval, the LLMs have already understood the tasks in shallower layers. 3) For Counterfact, StrategyQA, and Coinflip, The tasks are more difficult to understand, compared with others. Therefore, we consider the tasks in type 1 and 2 easy tasks, and those in type 3 are complex tasks.
5.2 Comparison Among the Number of Parameters
This section offers a comparative analysis of LLMs within their respective families, examining both accuracy levels and converging points across the models. Figure 3 reveals two recurring patterns within these families: for the tasks whose accuracy improved significantly by model learning, larger models tend to show converging points at earlier layers, suggesting they achieve their own peak comprehensions of concepts at lower layers; for the tasks whose accuracy do not change a lot, all LLM models show the converging points at early layers.
Two notable exceptions to this trend appear in the QWen family, over the Coinflip and IMDb datasets. For Coinflip, larger models exhibit delayed convergence. This deviation underscores the complexity of the reasoning required, illustrating how this task challenges even the larger models to extend their depth of understanding further. For IMDb, converging points fluctuate with the increasing size of the model because the number of layers is different among different sizes of LLMs, which amplifies the differences. These exceptions are also found in the Gemma family.
Furthermore, in Figure 3, we explore the peak accuracy levels across all layers for LLMs of differing sizes. The overarching trend indicates that larger models consistently achieve superior peak performance. This observation not only supports the notion that scaling up models enhances their effectiveness but also suggests that larger models develop more robust internal representations, validating the benefits of training models with greater capacity.
By comparing different sizes of models from the same LLM family, we have two observations. 1) As the number of parameters increases, peak accuracy gradually increases, and the converging point gradually advances. 2) Larger models grasp the concepts earlier and better.
5.3 Comparison Among the LLM Families
In this section, we examine how LLMs from various families, possessing a similar parameter count, process concepts as reflected by their converging points and peak accuracies. The overarching trends are highlighted in Figure 6, with detailed statistics on a layer-by-layer basis provided in Figure 5. Our findings reveal that while LLMs across different families may reach nearly identical peak accuracies, the layers at which they converge to these peaks can vary. For instance, in the HateEval and Counterfact datasets, we observe models converging at significantly deeper layers. This variation suggests that despite similar parameter scales, different models may employ varied mechanisms to tackle the same problems, reflecting the diversity in how models interpret and process complex information.
With the same magnitude of model parameters, the datasets generally perform comparably. The exceptions are Coinflip and STSA.
Figure 7: Noise and precision evaluation on STSA and Sarcasm of Gemma-2b.
5.4 Ablation Study
In this section, we investigate the impact of noise and precision reduction on the learning classification task of LLMs. For noise perturbation, we randomly add two short strings to the beginning of each problem and observe how the LLM’s performance changes. We also evaluate language models for 8, 16, and full 32 bits for precision. For example, Figure 7 illustrates the effect of adding noise and reducing bit representations to the Gemma-2b model. The addition of noise causes the learning curve to shift to the right, indicating a reduction in the speed of convergence. This suggests that the presence of noise in the input data can negatively impact the LLM’s learning process, slowing down its ability to converge to an optimal solution. We also see that there isn’t much difference between 32 and 16 bits, and the convergence rate becomes slower when we reduce the model to 8 bits. Therefore, we can compress the LLMs to 16 bits for future designs.
6 Discussion
Probing technology has the potential to revolutionize various aspects of machine learning and accelerate inference in large models. One exciting application involves using probes to determine the layer at which a concept is learned within a neural network in a large language model. This information can then be leveraged to perform targeted model pruning, specifically removing the layers that follow the identified layer. This pruning process can significantly speed up the inference process in large language models.
By deploying probes throughout a neural network, we can gather insights into how information is processed and transformed as it flows through the network. These probes act as detectors, identifying the layers where specific concepts or features are learned. This knowledge allows us to identify the critical layers that contribute significantly to the model’s final predictions.
Once the layer responsible for learning a particular concept is identified, targeted model pruning can be applied. Pruning involves removing unnecessary or less influential layers from the network while preserving its overall performance. By cutting off the layers that follow the identified layer, we eliminate computational overhead and reduce the model’s complexity, resulting in faster inference times.
7 Conclusions
This paper proposes the Concept Depth, the phenomenon that different concepts are learned in different layers of large language models, i.e. more difficult concepts are fully acquired with deeper layers. We conducted a number of experiments around Concept Depth using probe techniques. Our research suggests that LLMs tend to effectively categorize easy tasks, indicating that these concepts are learned in the first few initial layers. In contrast, complex tasks may only be recognizable (if at all) in deeper layers, and LLMs of the same size perform largely consistently across datasets in terms of concept depth.
Appendix A Appendix
Here, we provide our supplementary materials.
A.1 Metrics for Parts of the Layers
Table 3 shows the experimental results for the accuracy, F1-score, and AUC metrics of parts of the first, 25% depth, 50% depth, 67% depth, 83% depth, and the last layer of each model over the nine datasets.
Table 3: The experimental results of the nine datasets on nine LLMs. For each LLM, we select six layers (first, 25% depth, 50% depth, 67% depth, 83% depth, last) to record the accuracy, F1-score, and AUC.
Gemma-2b (18 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.556 0.601 0.588 0.635 0.667 0.626 0.446 0.411 0.422 0.582 0.61 0.584 0.502 0.509 0.518 0.74 0.822 0.742 0.666 0.729 0.68 0.722 0.788 0.725 0.731 0.808 0.714
25%-layer 0.602 0.642 0.629 0.62 0.64 0.6 0.94 0.987 0.939 0.637 0.683 0.633 0.675 0.745 0.664 0.793 0.89 0.794 0.872 0.942 0.874 0.884 0.953 0.885 0.857 0.935 0.852
50%-layer 0.639 0.7 0.665 0.65 0.705 0.632 0.983 0.999 0.983 0.648 0.699 0.642 0.729 0.826 0.716 0.802 0.891 0.8 0.915 0.972 0.917 0.941 0.982 0.941 0.89 0.955 0.887
67%-layer 0.683 0.751 0.708 0.695 0.783 0.69 0.992 1.0 0.992 0.671 0.727 0.666 0.766 0.852 0.758 0.809 0.891 0.806 0.929 0.976 0.93 0.936 0.984 0.936 0.893 0.96 0.889
83%-layer 0.62 0.668 0.637 0.585 0.625 0.579 0.985 0.999 0.985 0.652 0.703 0.645 0.704 0.782 0.696 0.807 0.89 0.803 0.911 0.97 0.914 0.926 0.977 0.926 0.882 0.944 0.878
last-layer 0.592 0.602 0.626 0.525 0.554 0.532 0.988 0.999 0.988 0.647 0.697 0.644 0.685 0.752 0.679 0.803 0.892 0.801 0.895 0.957 0.898 0.941 0.981 0.941 0.837 0.92 0.831
Gemma-7b (28 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.519 0.565 0.569 0.69 0.712 0.656 0.59 0.444 0.439 0.737 0.609 0.053 0.496 0.506 0.514 0.737 0.817 0.741 0.668 0.737 0.679 0.73 0.811 0.735 0.723 0.793 0.713
25%-layer 0.617 0.653 0.648 0.605 0.654 0.591 0.95 0.987 0.951 0.719 0.687 0.422 0.696 0.773 0.686 0.795 0.875 0.794 0.89 0.958 0.893 0.9 0.958 0.899 0.885 0.953 0.882
50%-layer 0.658 0.716 0.678 0.765 0.844 0.761 1.0 0.999 0.998 0.75 0.747 0.498 0.835 0.912 0.829 0.784 0.871 0.782 0.929 0.979 0.93 0.932 0.979 0.932 0.907 0.968 0.904
67%-layer 0.733 0.809 0.744 0.88 0.922 0.875 1.0 0.998 0.995 0.756 0.75 0.515 0.814 0.9 0.809 0.784 0.868 0.782 0.936 0.981 0.937 0.922 0.979 0.923 0.913 0.972 0.91
83%-layer 0.675 0.746 0.695 0.785 0.818 0.786 0.98 0.997 0.982 0.743 0.74 0.492 0.776 0.867 0.768 0.812 0.897 0.809 0.935 0.979 0.937 0.925 0.977 0.926 0.901 0.963 0.898
last-layer 0.604 0.666 0.625 0.57 0.602 0.578 0.95 0.996 0.972 0.759 0.748 0.481 0.729 0.808 0.721 0.82 0.901 0.817 0.92 0.975 0.922 0.938 0.984 0.938 0.862 0.932 0.86
LlaMA-7b (32 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.584 0.608 0.641 0.545 0.525 0.674 0.487 0.472 0.472 0.74 0.617 0.031 0.493 0.491 0.522 0.725 0.814 0.732 0.678 0.744 0.703 0.73 0.799 0.736 0.705 0.773 0.697
25%-layer 0.657 0.712 0.688 0.615 0.612 0.621 0.93 0.978 0.929 0.736 0.699 0.441 0.684 0.754 0.681 0.806 0.887 0.792 0.513 0.913 0.676 0.914 0.972 0.914 0.89 0.961 0.886
50%-layer 0.74 0.827 0.754 0.915 0.977 0.907 0.997 1.0 0.997 0.753 0.738 0.477 0.797 0.894 0.793 0.805 0.883 0.791 0.847 0.954 0.866 0.941 0.984 0.941 0.922 0.976 0.919
67%-layer 0.729 0.805 0.744 0.9 0.966 0.89 0.995 1.0 0.995 0.744 0.729 0.466 0.775 0.872 0.77 0.795 0.88 0.779 0.896 0.957 0.902 0.939 0.985 0.939 0.92 0.973 0.918
83%-layer 0.699 0.774 0.72 0.88 0.961 0.871 0.993 1.0 0.993 0.734 0.719 0.464 0.744 0.832 0.735 0.798 0.887 0.785 0.913 0.967 0.915 0.939 0.984 0.939 0.908 0.966 0.905
last-layer 0.67 0.744 0.69 0.815 0.9 0.8 0.993 1.0 0.993 0.743 0.731 0.464 0.739 0.818 0.731 0.831 0.914 0.83 0.935 0.984 0.937 0.944 0.987 0.944 0.895 0.955 0.892
LlaMA-13B (40 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.567 0.601 0.628 0.475 0.509 0.575 0.481 0.463 0.47 0.741 0.62 0.034 0.486 0.485 0.526 0.719 0.807 0.727 0.697 0.769 0.709 0.732 0.795 0.736 0.692 0.756 0.688
25%-layer 0.676 0.732 0.701 0.53 0.59 0.515 0.985 0.999 0.985 0.733 0.716 0.457 0.763 0.859 0.758 0.832 0.914 0.829 0.93 0.98 0.931 0.942 0.983 0.943 0.915 0.972 0.913
50%-layer 0.763 0.844 0.771 0.825 0.886 0.819 0.993 1.0 0.993 0.758 0.751 0.515 0.812 0.897 0.809 0.839 0.92 0.836 0.939 0.984 0.94 0.945 0.984 0.945 0.936 0.983 0.934
67%-layer 0.716 0.806 0.729 0.795 0.882 0.794 0.993 1.0 0.993 0.751 0.745 0.499 0.776 0.866 0.772 0.838 0.919 0.834 0.938 0.984 0.939 0.94 0.987 0.94 0.924 0.978 0.921
83%-layer 0.71 0.795 0.719 0.7 0.797 0.703 0.99 1.0 0.99 0.741 0.731 0.487 0.768 0.856 0.762 0.832 0.912 0.829 0.937 0.983 0.938 0.941 0.985 0.942 0.922 0.974 0.919
last-layer 0.693 0.772 0.704 0.645 0.715 0.664 0.99 1.0 0.99 0.75 0.743 0.499 0.76 0.841 0.752 0.835 0.913 0.833 0.935 0.984 0.937 0.946 0.988 0.946 0.91 0.969 0.908
QWen-0.5B (24 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.557 0.578 0.607 0.535 0.649 0.657 0.482 0.46 0.464 0.735 0.622 0.11 0.499 0.503 0.527 0.759 0.851 0.764 0.73 0.801 0.733 0.764 0.837 0.762 0.729 0.799 0.72
25%-layer 0.583 0.63 0.62 0.545 0.582 0.508 0.731 0.797 0.722 0.732 0.651 0.257 0.52 0.524 0.523 0.785 0.864 0.783 0.751 0.829 0.756 0.804 0.887 0.806 0.811 0.895 0.799
50%-layer 0.619 0.686 0.649 0.62 0.652 0.596 0.935 0.979 0.935 0.744 0.695 0.379 0.68 0.754 0.676 0.793 0.88 0.792 0.846 0.921 0.848 0.884 0.949 0.883 0.838 0.92 0.831
67%-layer 0.644 0.688 0.673 0.585 0.617 0.561 0.933 0.982 0.934 0.742 0.705 0.375 0.668 0.74 0.665 0.789 0.874 0.786 0.868 0.946 0.87 0.894 0.956 0.893 0.827 0.911 0.821
83%-layer 0.583 0.61 0.612 0.62 0.668 0.612 0.923 0.971 0.924 0.746 0.706 0.364 0.604 0.657 0.6 0.791 0.867 0.791 0.85 0.927 0.854 0.866 0.941 0.865 0.824 0.902 0.819
last-layer 0.55 0.567 0.584 0.55 0.613 0.541 0.912 0.971 0.912 0.742 0.703 0.357 0.579 0.616 0.579 0.784 0.866 0.781 0.844 0.922 0.848 0.879 0.951 0.88 0.825 0.9 0.82
QWen-1.8B (24 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.57 0.6 0.63 0.49 0.634 0.648 0.482 0.458 0.464 0.739 0.619 0.071 0.516 0.514 0.539 0.724 0.819 0.732 0.693 0.762 0.703 0.718 0.784 0.72 0.721 0.796 0.713
25%-layer 0.607 0.638 0.643 0.58 0.59 0.584 0.583 0.626 0.582 0.736 0.658 0.317 0.521 0.541 0.525 0.809 0.882 0.807 0.775 0.844 0.781 0.81 0.899 0.81 0.833 0.909 0.825
50%-layer 0.658 0.726 0.676 0.595 0.655 0.58 0.975 0.997 0.975 0.741 0.708 0.419 0.688 0.767 0.683 0.808 0.89 0.807 0.895 0.961 0.897 0.914 0.974 0.915 0.87 0.947 0.864
67%-layer 0.664 0.733 0.685 0.695 0.759 0.655 0.977 0.996 0.977 0.741 0.717 0.423 0.695 0.776 0.689 0.809 0.886 0.804 0.893 0.963 0.895 0.904 0.968 0.905 0.865 0.943 0.859
83%-layer 0.631 0.666 0.649 0.7 0.747 0.674 0.972 0.995 0.972 0.735 0.706 0.419 0.657 0.734 0.651 0.791 0.877 0.788 0.89 0.956 0.893 0.891 0.957 0.893 0.835 0.922 0.829
last-layer 0.595 0.642 0.604 0.615 0.667 0.605 0.973 0.996 0.973 0.74 0.704 0.404 0.638 0.713 0.631 0.795 0.877 0.794 0.879 0.949 0.882 0.906 0.962 0.907 0.812 0.896 0.808
QWen-7B (40 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.54 0.559 0.602 0.545 0.597 0.588 0.437 0.403 0.418 0.735 0.635 0.169 0.475 0.466 0.498 0.785 0.867 0.786 0.753 0.826 0.757 0.782 0.858 0.785 0.771 0.856 0.765
25%-layer 0.59 0.625 0.631 0.62 0.635 0.631 0.806 0.89 0.803 0.721 0.657 0.348 0.556 0.57 0.553 0.798 0.879 0.795 0.782 0.869 0.787 0.825 0.916 0.827 0.852 0.924 0.848
50%-layer 0.705 0.782 0.724 0.66 0.719 0.667 0.985 0.998 0.985 0.744 0.733 0.452 0.731 0.825 0.722 0.802 0.885 0.795 0.912 0.969 0.914 0.929 0.975 0.929 0.882 0.956 0.878
67%-layer 0.702 0.773 0.722 0.635 0.719 0.64 0.983 0.997 0.983 0.746 0.728 0.465 0.721 0.801 0.716 0.793 0.88 0.789 0.904 0.965 0.907 0.915 0.968 0.913 0.89 0.955 0.888
83%-layer 0.678 0.751 0.703 0.685 0.762 0.693 0.978 0.996 0.978 0.747 0.724 0.45 0.68 0.747 0.667 0.795 0.881 0.79 0.899 0.961 0.901 0.916 0.965 0.915 0.872 0.941 0.868
last-layer 0.676 0.718 0.693 0.585 0.623 0.587 0.982 0.995 0.981 0.75 0.727 0.453 0.657 0.718 0.643 0.782 0.869 0.778 0.902 0.964 0.905 0.92 0.973 0.919 0.848 0.921 0.845
QWen-14B (32 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.582 0.608 0.624 0.555 0.599 0.594 0.436 0.399 0.429 0.738 0.627 0.205 0.487 0.481 0.491 0.77 0.864 0.772 0.749 0.827 0.754 0.768 0.864 0.768 0.785 0.866 0.773
25%-layer 0.605 0.645 0.644 0.555 0.609 0.566 0.858 0.93 0.855 0.713 0.667 0.362 0.558 0.601 0.555 0.8 0.879 0.796 0.818 0.898 0.821 0.861 0.936 0.864 0.876 0.946 0.871
50%-layer 0.695 0.775 0.709 0.69 0.731 0.69 0.99 0.999 0.99 0.744 0.742 0.483 0.771 0.861 0.762 0.812 0.897 0.807 0.918 0.972 0.919 0.946 0.983 0.946 0.901 0.966 0.897
67%-layer 0.734 0.82 0.749 0.685 0.753 0.693 0.992 0.997 0.992 0.762 0.754 0.5 0.769 0.848 0.765 0.802 0.889 0.798 0.914 0.974 0.916 0.93 0.978 0.93 0.898 0.963 0.895
83%-layer 0.737 0.814 0.751 0.68 0.74 0.677 0.992 0.996 0.992 0.751 0.732 0.485 0.726 0.808 0.719 0.806 0.886 0.805 0.91 0.972 0.912 0.925 0.974 0.925 0.892 0.957 0.887
last-layer 0.714 0.785 0.729 0.65 0.681 0.646 0.982 0.996 0.981 0.753 0.74 0.483 0.707 0.779 0.701 0.802 0.881 0.798 0.908 0.97 0.91 0.932 0.979 0.933 0.871 0.943 0.867
QWen-14B (40 Layers) StrategyQA Coin Cities Common Claim Counterfact HateEval STSA IMDb Sarcasm
Metrics ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1 ACC AUC F1
1st-layer 0.569 0.59 0.626 0.605 0.639 0.646 0.461 0.422 0.448 0.733 0.62 0.162 0.481 0.486 0.504 0.766 0.86 0.769 0.75 0.824 0.753 0.778 0.856 0.776 0.772 0.859 0.762
25%-layer 0.603 0.637 0.631 0.68 0.704 0.66 0.866 0.945 0.865 0.72 0.669 0.393 0.621 0.67 0.615 0.8 0.89 0.797 0.842 0.919 0.845 0.862 0.939 0.861 0.874 0.946 0.869
50%-layer 0.735 0.835 0.747 0.71 0.754 0.681 0.995 1.0 0.995 0.76 0.748 0.515 0.791 0.875 0.787 0.808 0.891 0.807 0.932 0.981 0.934 0.935 0.984 0.935 0.919 0.974 0.916
67%-layer 0.798 0.883 0.808 0.7 0.791 0.674 0.995 1.0 0.995 0.761 0.76 0.502 0.775 0.859 0.769 0.816 0.894 0.814 0.933 0.98 0.934 0.938 0.978 0.937 0.914 0.968 0.911
83%-layer 0.796 0.872 0.807 0.715 0.768 0.705 0.992 1.0 0.992 0.753 0.735 0.493 0.753 0.843 0.748 0.816 0.896 0.811 0.927 0.978 0.927 0.936 0.978 0.936 0.912 0.964 0.91
last-layer 0.783 0.863 0.792 0.605 0.67 0.586 0.99 1.0 0.99 0.757 0.749 0.481 0.725 0.814 0.723 0.815 0.897 0.812 0.922 0.978 0.923 0.931 0.978 0.93 0.89 0.952 0.887
A.2 Examples of Each Dataset
For the 9 bi-classification tasks, we provide two examples from each class for each task. The blue sentences are the data samples, while the black sentences are the prompts.
Cities
"True" example:
Judge the statement is True or False. The city of Tokyo is in Japan.
"False" example:
Judge the statement is True or False. The city of Lodz is in the Dominican Republic.
CommonClaim
"True" example:
Judge the statement is True or False. Salmon will often return to the same spawning ground where then were born.
"False" example:
Judge the statement is True or False. A chicken has two right wings.
Counterfact
"True" example:
Judge the statement is True or False. The city of Tokyo is in Japan.
"False" example:
Judge the statement is True or False. Kanata South Ward is in Wisconsin.
HateEval
"Yes" example:
Here it is not about Refugees or Illegal immigrants. It is about whether one has documents before 1971 or not. Now, it is difficult for slum people and beggars to show valid documents, except the name in voter list. According to the comment, tell whether they present hate speech or not.
"No" example:
Labor migrants transfer almost \$10 billion a year to Ukraine. According to the comment, tell whether they present hate speech or not.
STSA
"Positive" example:
The production values are of the highest and the performances attractive without being memorable. The sentence above is a movie review and reflects the writer’s overall intention for this review. According to the sentence, judge whether the emotion is Positive or Negative.
"Negative" example:
Less a story than an inexplicable nightmare, right down to the population’s shrugging acceptance to each new horror. The sentence above is a movie review and reflects the writer’s overall intention for this review. According to the sentence, judge whether the emotion is Positive or Negative.
IMDb
"Positive" example:
This is the definitive movie version of Hamlet. Branagh cuts nothing, but there are no wasted moments. According to the movie review, judge whether it is Positive or Negative.
"Negative" example:
This is without a doubt the worst movie I have ever seen. It is not funny. It is not interesting and should not have been made. According to the movie review, judge whether it is Positive or Negative.
Sarcasm
"Yes" example:
Task: Detect sarcasm, help me identify whether this sentence is sarcastic. First, we need to understand what sarcasm is. Sarcasm is a form of verbal irony, where the intended meaning of the words is the opposite of the literal meaning. In other words, the speaker is saying one thing but meaning the opposite. Bashar al-Assad tries a tiny bit of sarin gas on self to see what it’s like. Think carefully according to the sentence. Is there any sarcasm in this sentence? Please answer Yes or No.
"No" example:
Task: Detect sarcasm, help me identify whether this sentence is sarcastic. First, we need to understand what sarcasm is. Sarcasm is a form of verbal irony, where the intended meaning of the words is the opposite of the literal meaning. In other words, the speaker is saying one thing but meaning the opposite. This ceo will send your kids to school, if you work for his company. Think carefully according to the sentence. Is there any sarcasm in this sentence? Please answer Yes or No.
StrategyQA
Note: This dataset was created in 2021. Queen Elizabeth was alive then.
"Yes" example:
Judge the question is true or false? Q: Will Queen Elizabeth be buried in the Pantheon? Let us think step by step. The stem of the sentence is Queen Elizabeth, burial, pantheon. Inference: First, the Pantheon is a church, so it is possible that she could be buried there. Second, Queen Elizabeth II is still alive, so she has not been buried yet. Third, even if she were to be buried in the Pantheon, it is unlikely that we would know about it ahead of time, so it is hard to say for sure. pred_ans: no. Do hamsters provide food for any animals? Let us think step by step...
"No" example:
Judge the question is true or false? Q: Will Queen Elizabeth be buried in the Pantheon? Let us think step by step. The stem of the sentence is Queen Elizabeth, burial, pantheon. Inference: First, the Pantheon is a church, so it is possible that she could be buried there. Second, Queen Elizabeth II is still alive, so she has not been buried yet. Third, even if she were to be buried in the Pantheon, it is unlikely that we would know about it ahead of time, so it is hard to say for sure. pred_ans: no. Could a llama birth twice during the War in Vietnam (1945-46)? Let us think step by step...
Coinflip
"Yes" example:
A coin is heads up. Whitney flips the coin. Erika does not flip the coin. Tj does not flip the coin. Benito flips the coin. Is the coin still heads up? Note that "flip" here means "reverse". According to the flipping process above, determine if a coin remains heads up after it is either flipped or left unflipped by individuals. Therefore, the answer (Yes or No) is?
"No" example:
A coin is heads up. Lucky does not flip the coin. Mireya flips the coin. Jj flips the coin. Kc flips the coin. Is the coin still heads up? Note that "flip" here means "reverse". According to the flipping process above, determine if a coin remains heads up after it is either flipped or left unflipped by individuals. Therefore, the answer (Yes or No) is?
A.3 LLM structure
Here, we give an introduction to the model structure (using LLaMA2-7B as an example).
LlamaForCausalLM(
    (model): LlamaModel(
        (embed_tokens):Embedding(32000,4096)
            (layers):ModuleList(
                (0-31):32 x lamaDecoderLayer(
                    (self_attn): LlamaSdpaAttention(
                        (g_proj):Linear(in_features=4096,out_features=4096,bias=False)
                        (k proj): Linear(in_features=4096,out_features=4096,bias=False)
                        (v proj):Linear(in_features=4096,out_features=4096, bias=False)
                        (o proj):Linear(in_features=4096, out_features=4096, bias=False)
                        (rotary emb):LlamaRotaryEmbedding()
                    (mlp):LlamaMLP(
                        (gate_proj): Linear(in_features=4096,out_features=11008,bias=False)
                        (up_proj): Linear(in_features=4096,out_features=11008,bias=False)
                        (down_proj): Linear(in_features=11008, out_features=4096, bias=False)
                        (act_fn): siLu()
                    (input_layernorm):LlamaRMSNorm()
                    (post_attention_layernorm):LlamaRMSNorm()
            (norm):LlamaRMSNorm()
   (lm_head): Linear(in_features=4096,out_features=32000,bias=False)